Two balls A and B of masses 10g and 20g moves on x axis towards each other with equal speed. They collide and stick together. What percentage of kinetic energy is lost?


1
Expert's answer
2021-07-05T08:42:28-0400

"m_a= 10g=0.01kg"

"m_b= 20g=0.02kg"

"\vec v_a= -\vec v_b"

"\text{according to the law of conservation of momentum:}"

"m_a\vec v_a+m_b\vec v_b=(m_a+m_b)\vec v"

"-m_a\vec v_b+m_b\vec v_b=(m_a+m_b)\vec v"

"\vec v=\frac{m_b-m_a}{m_a+m_b}\vec v_b"

"\vec v=\frac{\vec v_b}{3}"

"\text{before collision}"

"E_k= \frac{m_av_a^2}{2}+\frac{m_bv_b^2}{2}"

"E_k= \frac{m_av_b^2}{2}+\frac{m_bv_b^2}{2}"

"E_k= \frac{v_b^2(m_a+m_b)}{2}=0.015v_b^2"

"\text{after collision:}"

"E'_k= \frac{v^2(m_a+m_b)}{2}=\frac{v_b^2(m_a+m_b)}{18}=\frac{0.01v_b^2}{6}"

"p = \frac{E_k-E'_k}{E_k}=1-\frac{E'_k}{E_k}"

"p = 1- \frac{0.01v_b^2}{6*0.015*v_b^2}\approx0.89=89 \%"

"\text{Answer: 89\% of kinetic energy is lost}"